OpenStudy (anonymous):

What is the slope of the curve: r=1+2cos(theta) @ (theta)=pi/2

OpenStudy (anonymous):

\[r=1+2\cos(\theta) ;(\theta)=\pi/2\]

OpenStudy (diamondboy):

first do d derivative

OpenStudy (diamondboy):

can u?

OpenStudy (anonymous):

dr/dtheta = theta -2sin(theta)

OpenStudy (anonymous):

\[dr/d \theta = \theta -2\sin(\theta)\]

OpenStudy (anonymous):

simply plug in pi/2 for theta now

OpenStudy (anonymous):

easy

OpenStudy (mathmale):

\[r=1+2\cos(\theta) ;(\theta)=\pi/2\]Find dr/d(theta). The correct result does not begin with theta.

OpenStudy (mathmale):

Hints: What is the derivative of a constant? What is the derivative of the sine function?